मराठी

Two Unbiased Dice Are Thrown. Find the Probability That: (I) Neither a Doublet Nor a Total of 8 Will Appear

Advertisements
Advertisements

प्रश्न

Two unbiased dice are thrown. Find the probability that  neither a doublet nor a total of 8 will appear

Advertisements

उत्तर

We know that in a single throw of two dices, the total number of possible outcomes is (6 × 6) = 36.
Let S be the sample space.
Then n(S) = 36

Let E1 = event of getting neither a doublet nor a total of 8
Then E1' = event of getting either a doublet or a total of 8
∴ E1' = {(1, 1), (2, 2), (3, 3), (4, 4), (5, 5), (6, 6), (2, 6), (3, 5), (5, 3), (6, 2)}
i.e. n(E1') = 10
Thus, P(E1') = \[\frac{10}{36}\]

Hence, required probability P(E1) = 1- P(E1')

\[= 1 - \frac{10}{36} = \frac{36 - 10}{36} = \frac{26}{36} = \frac{13}{18}\]
 

 

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 33: Probability - Exercise 33.3 [पृष्ठ ४७]

APPEARS IN

आर.डी. शर्मा Mathematics [English] Class 11
पाठ 33 Probability
Exercise 33.3 | Q 28.1 | पृष्ठ ४७

व्हिडिओ ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्‍न

Which of the following can not be valid assignment of probabilities for outcomes of sample space S = {ω1, ω2,ω3,ω4,ω5,ω6,ω7}

Assignment ω1 ω2 ω3 ω4 ω5 ω6 ω7
(a) 0.1 0.01 0.05 0.03 0.01 0.2 0.6
(b) `1/7` `1/7` `1/7` `1/7` `1/7` `1/7` `1/7`
(c) 0.1 0.2 0.3 0.4 0.5 0.6 0.7
(d) –0.1 0.2 0.3 0.4 -0.2 0.1 0.3
(e) `1/14` `2/14` `3/14` `4/14` `5/14` `6/14` `15/14`

A coin is tossed twice, what is the probability that at least one tail occurs?


A fair coin is tossed four times, and a person win Re 1 for each head and lose Rs 1.50 for each tail that turns up.

From the sample space calculate how many different amounts of money you can have after four tosses and the probability of having each of these amounts.


Three coins are tossed once. Find the probability of getting

  1. 3 heads
  2. 2 heads
  3. at least 2 heads
  4. at most 2 heads
  5. no head
  6. 3 tails
  7. exactly two tails
  8. no tail
  9. atmost two tails.

Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.35 ... 0.25 0.6

Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.5 0.35 .... 0.7

4 cards are drawn from a well-shuffled deck of 52 cards. What is the probability of obtaining 3 diamonds and one spade?


If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the digits are repeated?


Two unbiased dice are thrown. Find the probability that the sum of the numbers obtained on the two dice is neither a multiple of 2 nor a multiple of 3


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that  all the balls are of different colours.


A bag contains 5 red, 6 white and 7 black balls. Two balls are drawn at random. What is the probability that both balls are red or both are black?


If a letter is chosen at random from the English alphabet, find the probability that the letter is a consonant .


In a lottery, a person chooses six different numbers at random from 1 to 20, and if these six numbers match with six number already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game?


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(ii)
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]

In a single throw of three dice, find the probability of getting the same number on all the three dice.


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that: all 10 are defective


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability thatat least one is defective


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is blue or white


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is red or yellow and numbered 1, 2, 3 or 4


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is white and numbered higher than 12 or yellow and numbered higher than 26.


Three-digit numbers are formed using the digits 0, 2, 4, 6, 8. A number is chosen at random out of these numbers. What is the probability that this number has the same digits?


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that C will be selected?


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are white


If the letters of the word ASSASSINATION are arranged at random. Find the probability that four S’s come consecutively in the word


If the letters of the word ASSASSINATION are arranged at random. Find the probability that two I’s and two N’s come together


If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together


If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together


Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.


While shuffling a pack of 52 playing cards, 2 are accidentally dropped. Find the probability that the missing cards to be of different colours ______.


The probability that a person visiting a zoo will see the giraffee is 0.72, the probability that he will see the bears is 0.84 and the probability that he will see both is 0.52.


The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get compartment is 0.96.


The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×